### v1

@@ Abstract
Context:
Aims:Because of their limited angular resolution, far-infrared telescopes are usually affected by confusion phenomenon. Since several galaxies can be located in the same instrumental beam, only the brightest objects emerge from the fluctuations caused by fainter sources. The probe far-infrared mission for astrophysics imager (PRIMAger) will observe the mid- and far-infrared (25–235 m) sky both in intensity and polarization. We aim to provide predictions of the confusion level and its consequences for future surveys.
Methods:We produced simulated PRIMAger maps affected only by the confusion noise using the simulated infrared extragalactic sky (SIDES) semi-empirical simulation. We then estimated the confusion limit in these maps and extracted the sources using a basic blind extractor. By comparing the input galaxy catalog and the extracted source catalog, we derived various performance metrics as completeness, purity, and the accuracy of various measurements (e.g., the flux density in intensity and polarization or the polarization angle).
Results:In intensity, we predict that the confusion limit increases rapidly with increasing wavelength (from 21 Jy at 25 m to 46 mJy at 235 m). The confusion limit in polarization is more than two orders of magnitude lower (from 0.03 mJy at 96 m to 0.25 mJy at 235 m). The measured flux density is dominated by the brightest galaxy in the beam, but other objects also contribute at longer wavelength (30 % at 235 m). We also show that galaxy clustering has a mild impact on confusion in intensity (up to 25 %), while it is negligible in polarization. In intensity, a basic blind extraction will be sufficient to detect galaxies at the knee of the luminosity function up to z3 and 10 M main-sequence galaxies up to z5. In polarization for the most conservative sensitivity forecast, 200 galaxies can be detected up to z=1.5 in two 1 500 h surveys covering 1 deg and 10 deg. For the baseline sensitivity, we expect 8 000 detections up to z=2.5 opening a totally new window on the high-z dust polarization. Finally, we show that intensity surveys at short wavelength and polarization surveys at long wavelength tend to reach confusion at similar depth. There is thus a strong synergy between them.
Conclusions:
Confusion of extragalactic sources in the far infrared: a baseline assessment of the performance of PRIMAger in intensity and polarization
@@ Conclusions:
The PRobe far-Infrared Mission for Astrophysics (PRIMA) project uses a 1.8 m space-based telescope cryogenically-cooled to 4.5 K with new generation detectors that take full advantage of the low thermal background. One of the two payload instruments is the PRIMA imaging camera (PRIMAger, Burgarella et al. 2023; Meixner et al. 2024). PRIMAger has two main bands. The first one is an hyperspectral band (PRIMAger Hyperspectral Imaging: PHI) which provides imaging with a linear variable filter at a spectral resolution R10 over 25 to 80 m. The second band (PRIMAger Polarization Imaging: PPI) provides 4 broad band filters between 91 and 235 m, sensitive to polarization. PRIMAger will operate with 100 mK cooled kinetic inductance detectors which development allows for an incomparable improvement of sensitivity in far infrared.
@@
An observatory like PRIMA will cover a wide range of science topics such as, but not limited to, origins of planetary atmospheres, evolution of galaxies, and build-up of dust and metals through cosmic time (Moullet et al., 2024). In addition, it will offer for the first time spaceborne high-sensitivity far-infrared polarimetric capabilities in the far infrared. So far no high-z source has been polarimetrically detected at these wavelengths, and only a strongly-lensed starburst galaxy at z=2.6 has been detected in the sub-millimeter (Geach et al., 2023).
@@
The goal of this paper is to assess the impact of the confusion on PRIMAger performance both in intensity and polarization based on the simulated infrared dusty extragalactic sky (SIDES, Béthermin et al. 2022) tools, and demonstrate the feasibility of high-z polarization surveys. It is important to realize that confusion arising from simple source extractions, as assessed here, is not an ultimate limit for well designed surveys and instrumentation. In this paper, we will focus on the ”classical” confusion limit for basic blind source extractors. Super-resolution techniques can certainly break through the classical confusion limit determined in these calculations to extract accurate SED information from fainter sources. More importantly, prior information derived from catalogs at un-confused wavelengths is very effective at improving the flux extraction to much fainter limits. PRIMAger’s hyperspectral architecture is especially designed to take advantage of this type of technique to break through confusion. In this paper we focus on performances expected from basic source extractors, while prior-based extractors, which show improvements by factors of several beyond the confusion noise established here, will be discussed in a companion paper (Donnellan et al., 2024).
@@ 2.1 The SIDES simulation
The confusion phenomena is highly connected to the flux density distribution of galaxies (e.g., Condon, 1974; Dole et al., 2004) and mildly by their spatial distribution as we will show in this study. To produce accurate forecasts of the confusion limit in the far infrared, we thus need a realistic model of the statistical source properties at these wavelengths.
The simulated infrared dusty extragalactic sky (SIDES, Béthermin et al. 2017111The material associated to SIDES can be found at https://data.lam.fr/sides/home.) is a semi-empirical model populating dark-matter halos from numerical simulations using recent observed physical relations. In this paper, we use the 2 deg version of SIDES. It connects the halo mass to the stellar mass using a sub-halo abundance matching technique (e.g., Behroozi et al., 2013). A fraction of the galaxies are drawn to be star-forming based on their stellar mass and redshift, and emits in the far-infrared. Their star formation rate (SFR) is then drawn based on the evolution of the main-sequence of star-forming galaxies, which is the relation between SFR and the stellar mass evolving with redshift (e.g, Schreiber et al., 2015). The observed scatter around this relation is taken into account by SIDES, and a population of high-SFR outliers is labeled as starbursts. Different spectral energy distributions (SEDs) are then attributed to galaxies depending if they are starbursts or not. These SEDs evolves with redshift following the observations of warmer dust at higher redshift (e.g., Béthermin et al., 2015). A temperature scatter on the SED templates is also included in the simulation. The AGN contribution is not included in the model, but Béthermin et al. (2012a) showed that it has a small impact on the galaxy number counts and thus the confusion noise (see Eq. 2 for the link between number counts and confusion noise).
@@
Finally, the CIB anisotropies (e.g., Planck Collaboration et al. (2014); Viero et al. (2013)), which is the integrated background from dust emission of galaxies at all redshifts, are also correctly recovered by SIDES, including cross-power spectra between different wavelengths (Béthermin et al., 2017; Gkogkou et al., 2023). The CIB anisotropies at small scale are dominated by the shot noise from galaxies below the detection threshold, while the signal at large scale is caused by the galaxy clustering. This agreement enhances our confidence that our model characterises both the clustering and flux distribution of faint sources. The faithfully reproduced cross-power spectra indicate that the galaxy colors are also reasonable.
@@ 2.2 Simulated PRIMAger maps
The confusion limit is the faintest flux density at which we can extract sources reliably in the limit of zero instrumental noise. In this paper, the instrumental noise refers both to the noise coming from the instrument itself and the photon noise from the various diffuse astrophysical backgrounds. To estimate the confusion limit, we produced simulated PRIMAger maps without instrumental noise using SIDES. Our simulations do not contain Galactic cirrus, which can potentially impact the photometry but are usually very faint in fields chosen for the deep surveys.
@@ 2.3 Adding integrated polarization to SIDES
In SIDES, we draw the galaxy integrated polarization fractions () from a Gaussian model of the distribution observed in the local Universe by Lopez-Rodriguez et al. (2022) centered on 1 % and with a standard deviation of 0.3 %. The method used to derive these values is presented in Appendix 15. In the current version of the model, we do not consider a dependence of the polarization with wavelength or the presence of a starburst in the galaxy. Lopez-Rodriguez (2023) showed that starburst outflows can produce a specific wavelength-dependent signature. However, SIDES does not contain a model for outflow, and we could not calibrated an empirical dependence on these parameters using the SALSA sample since we did not find any statistically-significant effect on the integrated polarization (see Appendix 15).
As in Lagache et al. (2020), we neglect the intrinsic alignments between the integrated polarization angle of galaxies. This is primarily motivated by the small (5 %) probability of a spiral galaxy to be aligned with its neighbors found both in observations (Singh et al., 2015) and in simulations (Codis et al., 2018). We thus drew randomly the polarization angle from a uniform distribution between 0 and . The various integrated Stokes parameter (, ) and the polarized flux density () are then derived for each simulated galaxy using:
@@ 3.2 Background
This is not trivial for data from far-infrared observatories, since they are not usually absolute photometers. Real maps often have a zero mean enforced in order to filter out instrumental and celestial backgrounds and foregrounds. In addition, because of confusion, no regions is free from sources to be suitable for estimating the background. To illustrate this, we show the flux density distribution from our PRIMAger simulated maps in Fig. 2. These simulated maps have a true zero, which corresponds to the absence of any galaxy emission. However, diffuse background (e.g., cosmic microwave background) and foregrounds (e.g., zodiacal light) are not included in the simulation and nor do they include instrumental backgrounds. In practice, we cannot use this zero for the photometry, since it cannot be measured in real data. Furthermore the faint sources provide an unresolved background that we should remove for unbiased photometry.
@@ 3.3 Classical confusion limit estimates and source extraction threshold
where is the beam function, is the flux limit above which sources can be detected, the number of galaxies per flux interval and per solid angle (usually called differential number counts). This equation is implicit, since is usually defined to be in the confusion limited case and needs to be solved e.g. by iteration. The limit is essential as without the integral would be divergent. However, the choice of where to place the limit is inherently subjective and the choice of is a convention without a strong rationale222The commonly presented rationale is that sources above the limit can be identified and removed and so would not contribute to the residual fluctuations, but a threshold will correspond to different detection probabilities depending on the underlying source counts.
To extract sources from the simulated maps using the method described in Sect. 3.1, we set the detection threshold to after subtracting the mode. The surface density of the detected sources goes from 43 962 sources per deg in the PHI1_1 band to 306 sources per deg in the PPI4 band. All the values are provided in Table 2. The photometry is also performed on the mode-subtracted map.
@@ 3.5 Completeness and purity estimates
The completeness as a function of the flux density is a key characterization of the source detection performance. A classical definition of the completeness is the fraction of sources at a given flux density in the input catalog (in our analysis, the SIDES simulated galaxy catalog) that are recovered in the output catalog produced by the source extractor. However, since several simulated galaxies can be found in the beam of a single source, recovery can be ambiguous. In this paper, we will use two definitions of a recovered source. In the definition A, we consider that a galaxy from the simulated catalog is recovered if it is located in the half-light radius of any source extracted from the associated simulated map. However, in this case, the completeness does not tend to zero at low flux density. At first order (no clustering), it converges to the fraction of the map encircled in the half-light radii of the extracted sources (up to 7 %). To avoid this problem, we introduce a definition B, where the galaxy must satisfy the additional condition of being the brightest source in half-light radius. The results are discussed in Sect. 4.2 and 5.3.
@@ 4.1 Photometric accuracy
In Fig. 3, we show the ratio between the flux density of the brightest galaxy in the beam and the measurement in the map. This is the inverse of the classical output versus input ratio used to illustrate the flux boosting, but our choice has the advantage to provide immediately the relative contribution of the brightest galaxy. In PHI bands, the median ratio is above 0.91 at the classical confusion limit and converges rapidly to unity at higher flux density. We also looked at the distributions around the median using the 16–84 % and 2.3–97.7 % ranges, (corresponding to 1 and 2 for Gaussian distributions). We note that the distribution is highly asymmetrical. Only rare outliers have an underestimated flux density (ratio1), while the lowest 2.3 % can have a flux density underestimated by half in these bands. This is a consequence of the large tail of positive outliers in the histogram of pixel flux densities (see Fig. 2). These overestimated flux densities are mainly caused by the blending of two sources with a similar flux density. Advanced deblending algorithms are usually very efficient to mitigate this effect (Donnellan et al., 2024).
In PPI bands, the contribution from other sources in the beam becomes more significantly with larger beams at longer wavelength. The median flux density ratio at the classical confusion limit decreases with increasing wavelength from 0.90 to 0.72. This effect has already been discussed in the case of the Herschel/SPIRE instrument by Scudder et al. (2016) and Béthermin et al. (2017), and is mainly caused by sources at other redshifts while a small contribution (5 %) comes from physically-related sources. The dispersion of the ratio around the median value also becomes more symmetrical with increasing wavelength. This is expected since the histogram of pixel flux densities becomes more symmetrical at longer wavelength (see Fig. 2). This is consequence of the central limit theorem as the approximately Poisson distribution of source fluxes becomes more Gaussian with a larger number of source per beam.
@@
We also estimated the flux uncertainties using half of the 84–16 % interval (corresponding to for a Gaussian distribution). At the classical confusion limit, in PHI bands, the relative flux uncertainties range from 8 % in band PHI1_1 to 15 % in in band PHI2_6, and are better than the 20 % expected from our 5 construction of the extraction threshold. This is because the distribution is highly non-Gaussian and the clipped variance is more sensitive to outliers. These relative uncertainties below 20 % confirm that our estimate of classical 5 confusion level is conservative in these bands, and deeper catalogs can be obtained with more advanced extraction methods (see Donnellan et al. 2024). In PPI bands, the performance degrades with increasing wavelength (15 % in PPI1, 19 % in PPI2, 22 % in PPI3, 26 % in PPI4). The accuracy in PPI3 and PPI4 bands is slightly worse than the 20% expected. In these bands, blind-extracted catalogs will be challenging to use in intensity, and their interpretation will either require complex statistical corrections of the fluxes or incorporating the effects of angular resolution through statistical models (e.g., Hayward et al., 2013; Cowley et al., 2015; Béthermin et al., 2017; Bing et al., 2023). In contrast, techniques such as prior-based source extraction will enable accurate measures of the fluxes (e.g., Donnellan et al. 2024).

### v3

@@ Abstract
Aims. Because of their limited angular resolution, far-infrared telescopes are usually affected by confusion phenomenon. Since several galaxies can be located in the same instrumental beam, only the brightest objects emerge from the fluctuations caused by fainter sources. The probe far-infrared mission for astrophysics imager (PRIMAger) will observe the mid- and far-infrared (25–235 m) sky both in intensity and polarization. We aim to provide predictions of the confusion level and its consequences for future surveys.
Methods. We produced simulated PRIMAger maps affected only by the confusion noise using the simulated infrared extragalactic sky (SIDES) semi-empirical simulation. We then estimated the confusion limit in these maps and extracted the sources using a basic blind extractor. By comparing the input galaxy catalog and the extracted source catalog, we derived various performance metrics as completeness, purity, and the accuracy of various measurements (e.g., the flux density in intensity and polarization or the polarization angle).
Results. In intensity maps, we predict that the confusion limit increases rapidly with increasing wavelength (from 21 Jy at 25 m to 46 mJy at 235 m). The confusion limit in polarization maps is more than two orders of magnitude lower (from 0.03 mJy at 96 m to 0.25 mJy at 235 m). Both in intensity and polarization maps, the measured (polarized) flux density is dominated by the brightest galaxy in the beam, but other objects also contribute in intensity maps at longer wavelength (30 % at 235 m). We also show that galaxy clustering has a mild impact on confusion in intensity maps (up to 25 %), while it is negligible in polarization maps. In intensity maps, a basic blind extraction will be sufficient to detect galaxies at the knee of the luminosity function up to z3 and 1011 M⊙ main-sequence galaxies up to z5. In polarization for the most conservative sensitivity forecast (payload requirements), 200 galaxies can be detected up to z=1.5 in two 1 500 h surveys covering 1 deg2 and 10 deg2. For a conservative sensitivity estimate, we expect 8 000 detections up to z=2.5 opening a totally new window on the high-z dust polarization. Finally, we show that intensity surveys at short wavelength and polarization surveys at long wavelength tend to reach confusion at similar depth. There is thus a strong synergy between them.
@@
The PRobe far-Infrared Mission for Astrophysics (PRIMA) project uses a 1.8 m space-based telescope cryogenically-cooled to 4.5 K with new generation detectors that take full advantage of the low thermal background. One of the two payload instruments is the PRIMA imaging camera (PRIMAger, Burgarella et al. 2023; Meixner et al. 2024). PRIMAger has two main bands. The first one is an hyperspectral band (PRIMAger Hyperspectral Imaging: PHI) which provides imaging with a linear variable filter at a spectral resolution R10 over 25 to 80 m. The second band (PRIMAger Polarization Imaging: PPI) provides 4 broad band filters between 91 and 235 m, sensitive to polarization. PRIMAger will operate with 100 mK cooled kinetic inductance detectors which allows for an incomparable improvement of sensitivity in far infrared.
@@
An observatory like PRIMA will cover a wide range of science topics such as, but not limited to, origins of planetary atmospheres, evolution of galaxies, and build-up of dust and metals through cosmic time (Moullet et al., 2024). In addition, it will offer for the first time spaceborne high-sensitivity far-infrared polarimetric capabilities in the far infrared. So far no high-z source has been polarimetrically detected at these wavelengths, and only one strongly-lensed starburst galaxy at z=2.6 has been published in the sub-millimeter (Geach et al., 2023). In addition, Chen et al. (2024) reported in a preprint a second sub-millimeter object at z=5.6 exhibiting kiloparsec-scale ordered magnetic fields while we were revising this paper.
Since high-z galaxies will not be spatially resolved, we will only detect the integrated polarization, if the magnetic fields driving the dust polarization in the various regions of a galaxy are ordered and their polarized flux densities add up at least partially. Else, in the disordered case, the signals from the various regions cancel out, leading to a very small integrated polarization. The integrated polarization fraction can vary with various physical parameters such as, e.g., the intrinsic dust polarization, the geometry of the galaxy, or the depolarization caused by turbulence (see, e.g., Sect. 6.3 of André et al. 2019). Pioneering studies in the local Universe showed that two main mechanisms lead to organized polarization patterns and thus significant integrated polarized fractions in star-forming galaxies: organized magnetic fields in disk galaxies (Lopez-Rodriguez et al., 2022) and starburst-driven outflows (Lopez-Rodriguez, 2023). In addition, active galactic nuclei (AGN) can also exhibit high polarization fractions in the far infrared (e.g., Lopez-Rodriguez et al., 2018; Marin et al., 2020), and could also lead to polarized outflows.
The goal of this paper is to assess the impact of the confusion on PRIMAger performance both in intensity and polarization based on the simulated infrared dusty extragalactic sky (SIDES, Béthermin et al. 2022) tools, and demonstrate the feasibility of high-z polarization surveys.
@@
It is important to realize that confusion arising from simple source extractions, as assessed here, is not an ultimate limit for well designed surveys and instrumentation. In this paper, we will focus on the ”classical” confusion limit for basic blind source extractors. Super-resolution techniques can certainly break through the classical confusion limit determined in these calculations to extract accurate SED information from fainter sources. More importantly, prior information derived from catalogs at un-confused wavelengths is very effective at improving the flux extraction to much fainter limits. PRIMAger’s hyperspectral architecture is especially designed to take advantage of this type of technique to break through confusion. In this paper we focus on performances expected from basic source extractors, while prior-based extractors, which show improvements by factors of several beyond the confusion noise established here, will be discussed in a companion paper (Donnellan et al., 2024).
@@
In this paper, we will use the terminology ”in intensity” to described all the quantities associated to standard photometric surveys and usually derived from specific intensity maps (e.g., flux density of point sources). The term ”in polarization” will corresponds to quantities extracted from the polarization maps (Sect. 2.3) such as polarized flux density or the polarized fraction of the flux density (shortened hereafter by polarization fraction).
@@ 2.1 The SIDES simulation
The confusion phenomenon is highly connected to the flux density distribution of galaxies (e.g., Condon, 1974; Dole et al., 2004) and mildly by their spatial distribution as we will show in this study. To produce accurate forecasts of the confusion limit in the far infrared, we thus need a realistic model of the statistical source properties at these wavelengths.
The simulated infrared dusty extragalactic sky (SIDES, Béthermin et al. 2017111The material associated to SIDES can be found at https://data.lam.fr/sides/home.) is a semi-empirical model populating dark-matter halos from numerical simulations using recent observed physical relations. In this paper, we use the 2 deg2 version of SIDES. It connects the halo mass to the stellar mass using a sub-halo abundance matching technique (e.g., Behroozi et al., 2013). A fraction of the galaxies are drawn to be star-forming based on their stellar mass and redshift, and emits in the far-infrared. Their star formation rate (SFR) is then drawn based on the evolution of the main-sequence of star-forming galaxies, which is the relation between SFR and the stellar mass evolving with redshift (e.g, Schreiber et al., 2015). The observed scatter around this relation is taken into account by SIDES, and a population of high-SFR outliers is labeled as starbursts. Different spectral energy distributions (SEDs) are then attributed to galaxies depending on whether they are starbursts or not. These SEDs evolve with redshift following the observations of warmer dust at higher redshift (e.g., Béthermin et al., 2015). A temperature scatter on the SED templates is also included in the simulation. The AGN contribution is not included in the model, but Béthermin et al. (2012a) showed that it has a small impact on the galaxy number counts and thus the confusion noise (see Eq. 2 for the link between number counts and confusion noise).
@@
Finally, the CIB anisotropies (e.g., Planck Collaboration et al. (2014); Viero et al. (2013)), which is the integrated background from dust emission of galaxies at all redshifts, are also correctly recovered by SIDES, including cross-power spectra between different wavelengths (Béthermin et al., 2017; Gkogkou et al., 2023). The CIB anisotropies at small scale are dominated by the shot noise from galaxies below the detection threshold, while the signal at large scale is caused by the galaxy clustering. This agreement enhances our confidence that our model characterizes both the clustering and flux distribution of faint sources. The faithfully reproduced cross-power spectra indicate that the galaxy colors are also reasonable.
@@ 2.2 Simulated PRIMAger maps
The confusion limit is the faintest flux density at which we can extract sources reliably in the limit of zero instrumental noise. In this paper, the instrumental noise refers both to the noise coming from the instrument itself and the photon noise from the various diffuse astrophysical backgrounds. To estimate the confusion limit, we produced simulated PRIMAger maps without instrumental noise using SIDES. Our simulations do not contain Galactic cirrus, which can potentially impact the photometry but are usually very faint in fields chosen for the deep surveys. However, in polarization, Galactic cirrus are expected to have a typically 5 times higher polarization fraction than unresolved galaxies (5 % versus 1 %, see Planck Collaboration et al. 2020 and Sect. 2.3). The survey footprints will thus have to be chosen very carefully to minimize the cirrus contamination in polarization.
@@ 2.3 Adding integrated polarization to SIDES
In SIDES, we draw the galaxy integrated polarization fractions () from a Gaussian model of the distribution observed in the local Universe by Lopez-Rodriguez et al. (2022) centered on 1 % and with a standard deviation of 0.3 %. The method used to derive these values is presented in Appendix 15. Since the 3- confidence interval of the central value of the polarized fraction goes from 0.7 % to 1.3 %, we will also discuss scenarios with these extreme values to illustrate the uncertainties on our forecasts. For simplicity, we assume that the relative scatter () on the polarized fraction is constant (0.3) in all the scenarios to be able to apply a simple rescaling to the polarized maps. In the current version of the model, we do not consider a dependence of the polarization with wavelength or the presence of a starburst in the galaxy. Lopez-Rodriguez (2023) showed that starburst outflows can produce a specific wavelength-dependent signature. However, SIDES does not contain a model for outflow, and we could not calibrate an empirical dependence on these parameters using the SALSA sample since we did not find any statistically-significant effect on the integrated polarization (see Appendix 15).
As in Lagache et al. (2020), we neglect the intrinsic alignments between the integrated polarization angles of galaxies. This is primarily motivated by the small (5 %) probability of a spiral galaxy to be aligned with its neighbors found both in observations (Singh et al., 2015) and in simulations (Codis et al., 2018). We thus drew randomly the polarization angle from a uniform distribution between 0 and . The various integrated Stokes parameter (, ) and the polarized flux density () are then derived for each simulated galaxy using:
@@ 3.2 Background
This is not trivial for data from far-infrared observatories, since they are not usually absolute photometers. Real maps often have a zero mean enforced in order to filter out instrumental and celestial backgrounds and foregrounds. In addition, because of confusion, no region is free from sources to be suitable for estimating the background. To illustrate this, we show the flux density distribution from our PRIMAger simulated maps in Fig. 2. These simulated maps have a true zero, which corresponds to the absence of any galaxy emission. However, diffuse background (e.g., cosmic microwave background) and foregrounds (e.g., zodiacal light) are not included in the simulation and nor do they include instrumental backgrounds. In practice, we cannot use this zero for the photometry, since it cannot be measured in real data. Furthermore the faint sources provide an unresolved background that we should remove for unbiased photometry.
@@ 3.3 Classical confusion limit estimates and source extraction threshold
where is the beam function, is the flux limit above which sources can be detected, the number of galaxies per flux interval and per solid angle (usually called differential number counts). This equation is implicit, since is usually defined to be in the confusion limited case and needs to be solved e.g. by iteration. The limit is essential as without the integral would be divergent. However, the choice of where to place the limit is inherently subjective and the choice of is a convention without a strong rationale444The commonly presented rationale is that sources above the limit can be identified and removed and so would not contribute to the residual fluctuations, but a threshold will correspond to different detection probabilities depending on the underlying source counts.
To extract sources from the simulated maps using the method described in Sect. 3.1, we set the detection threshold to after subtracting the mode. The photometry is also performed on the mode-subtracted map. The surface density of the detected sources goes from 43 962 sources per deg2 in the PHI1_1 band to 306 sources per deg2 in the PPI4 band. All the values are provided in Table 2. The surface density of sources above the confusion limit is slightly higher in polarization than in intensity in the same band (see discussion in Sect. 5.3). The confusion metrics expected for a different mean polarization fraction can be computed applying a (/1 %) scaling factor (except for the surface densities, see Appendix C).
@@ 3.5 Completeness and purity estimates
The completeness as a function of the flux density (polarized or not, in this section we use the term flux density to discuss both the intensity and polarization cases) is a key characterization of the source detection performance. A classical definition of the completeness is the fraction of sources at a given flux density in the input catalog (in our analysis, the SIDES simulated galaxy catalog) that are recovered in the output catalog produced by the source extractor. However, since several simulated galaxies can be found in the beam of a single source, recovery can be ambiguous. In this paper, we will use two definitions of a recovered source. In the definition A, we consider that a galaxy from the simulated catalog is recovered if it is located in the half-light radius of any source extracted from the associated simulated map. However, in this case, the completeness does not tend to zero at low flux density. At first order (no clustering), it converges to the fraction of the map encircled in the half-light radii of the extracted sources (up to 7 %). To avoid this problem, we introduce a definition B, where the galaxy must satisfy the additional condition of being the brightest source in half-light radius. The results are discussed in Sect. 4.2 and 5.3.
@@ 4.1 Photometric accuracy
In Fig. 3, we show the ratio between the flux density of the brightest galaxy in the beam and the measurement in the map. This is the inverse of the classical output versus input ratio used to illustrate the flux boosting, but our choice has the advantage to provide immediately the relative contribution of the brightest galaxy. In PHI bands, the median ratio is above 0.91 at the classical confusion limit and converges rapidly to unity at higher flux density. We also studied the distributions around the median using the 16–84 % and 2.3–97.7 % ranges (corresponding to 1 and 2 for Gaussian distributions). We note that the distribution is highly asymmetrical. Only rare outliers have an underestimated flux density (ratio1), while the lowest 2.3 % can have a flux density underestimated by half in these bands. This is a consequence of the large tail of positive outliers in the histogram of pixel flux densities (see Fig. 2). These overestimated flux densities are mainly caused by the blending of two sources with a similar flux density. Advanced deblending algorithms are usually very efficient to mitigate this effect (Donnellan et al., 2024).
In PPI bands, the contribution from other sources in the beam becomes more significant with larger beams at longer wavelength. The median flux density ratio at the classical confusion limit decreases with increasing wavelength from 0.90 to 0.72. This effect has already been discussed in the case of the Herschel/SPIRE instrument by Scudder et al. (2016) and Béthermin et al. (2017), and is mainly caused by sources at other redshifts while a small contribution (5 %) comes from physically-related sources. The dispersion of the ratio around the median value also becomes more symmetrical with increasing wavelength. This is expected since the histogram of pixel flux densities becomes more symmetrical at longer wavelength (see Fig. 2). This is consequence of the central limit theorem as the approximately Poisson distribution of source fluxes becomes more Gaussian with a larger number of source per beam.
@@
We also estimated the flux uncertainties using half of the 84–16 % interval (corresponding to for a Gaussian distribution). At the classical confusion limit, in PHI bands, the relative flux uncertainties range from 8 % in band PHI1_1 to 15 % in band PHI2_6, and are better than the 20 % expected from our 5 construction of the extraction threshold. This is because the distribution is highly non-Gaussian and the clipped variance is more sensitive to outliers. These relative uncertainties below 20 % confirm that our estimate of classical 5 confusion level is conservative in these bands, and deeper catalogs can be obtained with more advanced extraction methods (see Donnellan et al. 2024). In PPI bands, the performance degrades with increasing wavelength (15 % in PPI1, 19 % in PPI2, 22 % in PPI3, 26 % in PPI4). The accuracy in PPI3 and PPI4 bands is slightly worse than the 20% expected. In these bands, blind-extracted catalogs will be challenging to use in intensity, and their interpretation will either require complex statistical corrections of the fluxes or incorporating the effects of angular resolution through statistical models (e.g., Hayward et al., 2013; Cowley et al., 2015; Béthermin et al., 2017; Bing et al., 2023). In contrast, techniques such as prior-based source extraction will enable accurate measures of the fluxes (e.g., Donnellan et al. 2024).